Project

General

Profile

03042015

Attendance Ash River, Team IU, Jose, Leon, Louise

Maintenance was preformed on FD yesterday but not on ND.

Detectors are running very stable.

Can see increase number of channels on FD from work yesterday. Can also see a increase in the noise rate, perhaps larger than is explain by increased number of channels. Could be explained by DSO scan and natural fluctuations but we will keep an eye on it. Generally FD noise is reasonably constant.

The lights have been off on LL2 and LL3 all of last week and we have had not any large FEB drop outs. The week plots contain some information from some period of time with lights on, will be interesting to check tomorrow after a week has passed.

The near detector continues to get noisier, but do not see any effect on the physics only the noise slice. We can consider turning off the power if it gets to high or if we don't get an no-A147 for a while.
Leon points out that with the lower threshold that we run the ND at, the ND noise is much lower than for the FD.

  • Action: Leon asks for a toggle on/off button for the nearline plots. This is apparently already being worked on.

The tutorial videos went send out to the group. No comments expect that certain browsers seem to have issues playing them.

The new shift website is nearly ready to be sent to experts for comments and suggestions.